OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x^2, and g(x) = 108 when x = 6.

OpenStudy (campbell_st):

well let k be the constant of variation so the equation is \[g(x) = k \times x^2\] substitute x = 6 amg g(x) = 108 then solve for k hope it helps

OpenStudy (anonymous):

So g(6)=k*6^2 ?

OpenStudy (campbell_st):

not quite 108 = k * 6^2

OpenStudy (anonymous):

108=k*36 \[108/36 =k\]

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

So g(x)=3x^2?

OpenStudy (campbell_st):

thats correct